OpenStudy (anonymous):

my question is: 9x(5x2-8x+3) the 2 after the x is meant to be squared. I can't seem to figure it out.

OpenStudy (anonymous):

45x^3-72x^2+27x

OpenStudy (anonymous):

Distribute the 9x 9x(5x^2) = 45x^3 9x(8x) = 72x^2 9x(3) = 27x

OpenStudy (anonymous):

Thank you so much. :)

OpenStudy (anonymous):

40x2+40x+27 the 2 after x squared

OpenStudy (anonymous):

:( sorry put wrong number this is correct should be 45x2+40x+27